if the measure of an angle exceeds the measure of its compliment by 20, what is the measure of the angle?
Found 3 solutions by richwmiller, JimboP1977, Earlsdon:
Answer by richwmiller(17219)

They are just trying to confuse you with this fancy talk of exceed which just means it is more than.
The key question is what is the complement?
Complementary angles add up to 90.
so we have
x=y+20
x-y=20
x+y=90
add the last two equations
2x=110
x=55
y=90-55=35
check
35+55=90
55-35=20
ok

Answer by JimboP1977(311)

90 degrees minus the angle = complement of the angle.
Angle-20=Complement.
so 90-Angle=Angle-20
90=2*Angle-20
110=2*Angle
55=Angle.

Answer by Earlsdon(6294)

Let A be the angle, so its complement would be (90-A), and 20 more than that would be (90-A)+20.
So we can write an equation to find A.
A = (90-A)+20 Simplify.
A = 110-A Add A to both sides.
2A = 110 Divide both sides by 2.
A = 55
The measure of the angle is 55 degrees.